# Finding the unusual red giant remnants of cataclysmic variable mergers

### v1

@@ Abstract
Mergers between helium white dwarfs and main-sequence stars are likely common, producing red giant-like remnants making up roughly a few percent of all low-mass ($\lesssim 2M_{\odot}$) red giants. Through detailed modelling, we show that these merger remnants possess distinctive photometric, asteroseismic, and surface abundance signatures through which they may be identified. During hydrogen shell burning, merger remnants reach higher luminosities and possess pulsations which depart from the usual degenerate sequence on the asteroseismic $\Delta\nu$$\Delta\Pi$ diagram for red giant branch stars. For sufficiently massive helium white dwarfs, merger remnants undergo especially violent helium flashes which can dredge up a large amount of core material (up to $\sim 0.1M_{\odot}$) into the envelope. Such post-dredge-up remnants are more luminous than normal red clump stars, surface carbon-rich, and possess a wider range of asteroseismic g-mode period spacings and mixed-mode couplings. Recently discovered low-mass ($\lesssim 0.8M_{\odot}$) red clump stars may be core helium-burning remnants of mergers involving lower-mass helium white dwarfs.
@@ 1 Introduction
Growing evidence suggests that mergers between main-sequence (MS) stars and low-mass white dwarfs should be fairly common. The likely outcome of such a merger is to accrete the MS star onto the white dwarf, igniting hydrogen shell burning and creating an unusual red giant (RG) star. These RG remnants of those mergers should exist within the stellar population, and may be identifiable through a combination of their photometry, pulsations, and surface abundances.
Cataclysmic variables, or CVs, are stably mass-transferring systems with white dwarf (WD) accretors and MS donors. Typically, the progenitor binaries of CVs are post-common-envelope systems (Paczynski et al., 1976) which have been tightened by magnetic braking and gravitational waves until the onset of Roche lobe overflow (Knigge, 2011).
@@
    the typical accretor mass in CVs is $\simeq~{}0.8M_{\odot}$, greater than the typical mass of carbon–oxygen white dwarfs (CO WDs) $\simeq~{}0.6M_{\odot}$, and
@@
    no helium(-core) white dwarf (hereafter He WD) accretors have ever been discovered in CVs, even though they should be frequent outcomes of conventional binary stellar evolution (Zorotovic et al., 2011; Zorotovic & Schreiber, 2020).
@@
In the past decade, it was realized that the WD mass problem can naturally be solved by an extra, accretor mass-dependent angular momentum loss mechanism (Schreiber et al., 2015). This so-called consequential angular momentum loss would preferentially destabilize the mass transfer of CV-like systems with lower-mass white dwarfs, causing them to quickly merge before they can be observed (Belloni et al., 2018). Physically, this mechanism may be a frictional effect associated with nova events (Shen, 2015), which are expected to be much longer-duration for lower-mass white dwarfs (Shara et al., 1993), although the subsequent analysis is agnostic to the details.
@@
Figure 1: A summary diagram of the behavior of the He WD–MS merger remnants investigated in this work. On the RGB, merger remnants have cold and slightly more compact cores, which inflates their radii and changes the radiative structures of their cores to which asteroseismology is sensitive (Section 3). Sufficiently massive merger remnants ignite helium in abnormally energetic helium flashes. For massive enough progenitor He WDs, these events can dredge up significant amounts of helium into merger remnants’ envelopes (Section 4), changing their surface abundances, and greatly modifying both their structure and evolution. Less massive remnants become horizontal branch stars (Section 4.5), or even hot subdwarfs or He WDs (not explored in this work).
@@ 2 Stellar models
We do this in the following stages:
@@
    An isolated star with $M=1.2M_{\odot}$ is evolved through the MS and part of the RGB until it attains the desired initial core mass of the remnant, $M_{\mathrm{WD}}$. This initial mass is chosen to be close to typical RG masses from Kepler (e.g., Yu et al., 2018). Because RGs with $M\lesssim 2M_{\odot}$ all obey the same core mass–luminosity relation, the subsequent analysis is insensitive to this choice.
@@
    To stellar-engineer the merger remnant, we start with a scaffold RG model with core mass $M_{\mathrm{WD}}$, and replace shells with mass coordinates $m\leq M_{\mathrm{WD}}$ with the He WD model. To build the scaffold model, we start with the original RG model from above. We remove envelope mass from the scaffold model via relaxation with $|\dot{M}|=10^{-3}M_{\odot}\,\mathrm{yr}^{-1}$ with burning disabled, until the RG attains the desired final envelope mass $M_{\mathrm{MS}}$ (physically equal to the mass of the MS star participating in the merger, assuming no mass lost). This produces a merger remnant with total mass $M=M_{\mathrm{MS}}+M_{\mathrm{WD}}$. For most merger remnant models in this work, we arbitrarily fix $M_{\mathrm{MS}}$ such that $M=0.8M_{\odot}$, to resemble fairly typical low-mass RGs which may result from CV mergers.
@@
    To ensure numerical convergence, the resulting remnant is evolved for $2\,\mathrm{kyr}$ without burning and gold tolerances disabled. Nuclear reactions and gold tolerances are re-enabled at $2\,\mathrm{kyr}$ and $4\,\mathrm{kyr}$, respectively. Evolution through the helium flash involves disabling gold tolerances again, and the models are terminated after helium-burning when the central helium fraction drops to $Y_{c}\leq 10^{-3}$.
@@ 3.1 More compact cores inflate the radius
Figure 2: The trajectories on the Hertzsprung–Russell diagram of merger remnant models with varying He WD mass $M_{\mathrm{WD}}$ and total mass $M=0.80M_{\odot}$. The red dashed line indicates a non-merged model with $M=0.80M_{\odot}$. The top-left and bottom-left panels zoom into the tRGB and CHeB stages, respectively. The red-outlined star symbols on the top-left panel indicate the location of the helium flash. On the bottom-left panel, only evolution $\gtrsim 5\,\mathrm{Myr}$ following the helium flash is shown, and points are distributed $10\,\mathrm{Myr}$ apart. We have excised times near the helium flash, when the envelope sometimes becomes hydrodynamical.
@@ 3.2 Asteroseismic signatures on the red giant branch
Figure 3: The Brunt–Väisälä frequency $N$ (top) and temperature $T$ (bottom) shown for the $M_{\mathrm{MS}}=0.55M_{\odot}$, $M_{\mathrm{WD}}=0.25M_{\odot}$ merger remnant (solid gray-blue) and non-merged model of $M=0.80M_{\odot}$ (red dashed), at various stages of evolution. The two models are compared at equal core masses $M_{c}$. The shaded region denotes helium added to the helium core of the merger remnant model during its evolution.
@@
defined here specifically for the dipole ($\ell=1$) modes. This quantity probes the radiative core of the star. The integral in Equation 4 is taken over $\mathcal{R}$, the g-mode cavity of the RG, i.e., the region where the mode frequency $\omega<N$, where $N$ is the Brunt–Väisälä frequency. At present, $\Delta\Pi$ has been measured for a few thousand RGs (e.g., Vrard et al., 2016), observationally constraining the structures of their radiative cores.
@@
However, in a short time $<\tau_{\mathrm{th,core}}$, electronic thermal conduction (approximately obeying the nonlinear law in Equation B5) sets up a temperature gradient $\nabla$. Curiously, we find that this temperature gradient happens to mostly cancel out the factor $\sqrt{T_{\mathrm{shell}}/T}$ (see Equation 5) to cause the $N$ profiles of the single and merger remnant models to be similar (top center panels of Figure 3). Equality between the Brunt–Väisälä frequency profiles occurs long before the interior of the core has thermalized to the shell temperature. Therefore, in the relatively short time that it takes heat conduction to set up a temperature gradient throughout the core, the lower temperature of the merger remnant core no longer works to increase $\Delta\Pi$. The reason for this “coincidental” cancellation remains mysterious, but may be related to some properties of long-lived pre-thermalized solutions of nonlinear heat diffusion. These solutions (called “intermediate asymptotics”; Barenblatt, 1996) exhibit self-similar behavior in many nonlinear heat diffusion problems (often in the context of diffusion in a porous medium), and have been studied extensively in applied mathematics (by, e.g., Witelski & Bernoff, 1998; Galaktionov et al., 2004; Hayek, 2014).
@@
Figure 4: An asteroseismic spacing diagram ($\Delta\Pi$ versus $\Delta\nu$) showing merger remnant evolutionary sequences and a non-merged model (red dashed) all with $M=0.80M_{\odot}$. The top-right panel shows the entire parameter space, and the left and bottom panels zoom in on CHeB stars and the RGB, respectively. Roughly speaking, both RGB and CHeB stars evolve from right to left on this diagram (i.e., towards decreasing $\Delta\nu$, or increasing $R$). The black points are observations taken from the catalog of Gaulme et al. (2020). In the top-right panel, colored triangles denote the last $5\,\mathrm{Myr}$ of the CHeB phase.
@@
Much of the phase where merger remnants depart from the degenerate sequence takes place during the contraction phase soon after the formation of the remnant. This is only observable for remnants sufficiently low on the RGB where g modes are detectable (i.e., below the RGB bump, which is the “hook” feature in Figure 4). Hence, this sort of asteroseismic diagnosis may only be possible for sufficiently recent mergers involving sufficiently low-mass He WDs.
@@ 3.3 Overbright tip of the RGB
Figure 5: Top: The maximum helium-burning luminosities $L_{\mathrm{He,flash}}$ attained during the helium flash in our merger remnant models with total mass $M=M_{\mathrm{MS}}+M_{\mathrm{WD}}=0.80M_{\odot}$. Bottom: For the merger remnant models, we show core masses $M_{\mathrm{c@flash}}$ at the helium flash (purple circles), mass coordinates $M_{\mathrm{flash}}$ where the helium flash begins (magenta stars), and initial (turquoise triangles) and final (green squares) core masses $M_{\mathrm{CHeB,ci}}$ and $M_{\mathrm{CHeB,cf}}$ during the CHeB phase. In both panels, dashed lines denote values for the $M=0.80M_{\odot}$ non-merged model.
@@ 4 Core helium-burning phase
In our models, the maximum helium-burning luminosity attained by a merger remnant is roughly a linear function of $M_{\mathrm{WD}}$, with the most extreme model attaining a factor of $\approx\!5$ higher helium-burning luminosity than attained by a non-merged star (see the top panel of Figure 5). The energy production rate can be a whopping $\approx\!10^{10}L_{\odot}$ at the peak of the helium flash. Since mergers involving more massive WDs than those on our model grid are possible, even larger helium-burning luminosities may occur in nature. The most vigorous helium flash in our merger models also burns a larger fraction of the core into carbon, up to $\simeq 7\%$ by mass.
@@ 4.3 Asteroseismic mixed-mode coupling
Figure 10: Selected surface abundances $\mathrm{A}(\mathrm{X})$ (defined in Equation 11) during the CHeB phase. Red points denote the $M=0.80M_{\odot}$ non-merged model, and starred points denote post-dredge-up merger remnants. For ${}^{7}\mathrm{Li}$, filled symbols denote lithium enrichment due to the helium flash only (described in Section 4.4), whereas hollow symbols denote lithium enrichment due to both the helium flash and merger itself (which may be inaccurately modelled).
@@ 4.4 Abundance anomalies
We find that post-dredge-up remnants exhibit significant 12C surface enrichment, up to $\simeq 1\,\mathrm{dex}$. This enhancement arises from dredge-up of core material which has been partially fused during the helium flash into carbon. Post-dredge-up remnants thus possess increased values of $\mathrm{C}/\mathrm{Fe}$ and decreased values of ${}^{13}\mathrm{C}/{}^{12}\mathrm{C}$. Additionally, our post-dredge-up remnant models also possess significant surface enhancements in ${}^{18}\mathrm{O}$ and ${}^{22}\mathrm{Ne}$, which are created by successive $\alpha$ captures of ${}^{14}\mathrm{N}$ (e.g., Clayton, 2003) during the helium flash. While not included in our reaction network, it is also probable that a significant amount of ${}^{26}\mathrm{Mg}$ is formed through $\alpha$ capture of ${}^{22}\mathrm{Ne}$ in the abnormally hot helium flash (as in Shen et al., 2023). The surface abundance of ${}^{26}\mathrm{Mg}$ is therefore also likely to be enhanced.
The lithium abundance is also important, because it can be created by the burning of 3He or destroyed by burning at temperatures comparable to those required for hydrogen burning. Lithium-rich giants have previously been suggested to have been formed via binary interactions or mergers (e.g., Zhang & Jeffery, 2013; Casey et al., 2019). Our merger remnant models become highly lithium-rich very soon after merger, owing to a brief dredge-up event which occurs when hydrogen burning is first turned on (similar to the Cameron–Fowler mechanism; Cameron & Fowler, 1971). While such a merger-era lithium enhancement seems plausible (see, e.g., observations of lithium enrichment in luminous red novae; Kamiński et al., 2023), our treatment of the merger process itself is highly artificial, and the evolution of the remnant is unreliable for post-merger ages less than a thermal time, $\tau_{\mathrm{th,env}}$. Modification to the surface abundances of other species during this (possibly artificial) dredge-up event are minor, since they primordially occur in much higher abundance than lithium.

### v2

@@ Abstract
Mergers between helium white dwarfs and main-sequence stars are likely common, producing red giant-like remnants making up roughly a few percent of all low-mass ($\lesssim 2M_{\odot}$) red giants. Through detailed modeling, we show that these merger remnants possess distinctive photometric, asteroseismic, and surface abundance signatures through which they may be identified. During hydrogen shell burning, merger remnants reach higher luminosities and possess pulsations which depart from the usual degenerate sequence on the asteroseismic $\Delta\nu$$\Delta\Pi$ diagram for red giant branch stars. For sufficiently massive helium white dwarfs, merger remnants undergo especially violent helium flashes which can dredge up a large amount of core material (up to $\sim 0.1M_{\odot}$) into the envelope. Such post-dredge-up remnants are more luminous than normal red clump stars, are surface carbon-, helium-, and possibly lithium-rich, and possess a wider range of asteroseismic g-mode period spacings and mixed-mode couplings. Recent asteroseismically determined low-mass ($\lesssim 0.8M_{\odot}$) red clump stars may be core helium-burning remnants of mergers involving lower-mass helium white dwarfs.
keywords:
@@ 1 Introduction
Growing evidence suggests that mergers between main-sequence (MS) stars and low-mass white dwarfs should be fairly common (Schreiber et al., 2015). The likely outcome of such a merger is to accrete the MS star onto the white dwarf, igniting hydrogen shell burning and creating an unusual red giant (RG) star. These RG remnants of those mergers should exist within the stellar population, and may be identifiable through a combination of their photometry, pulsations, and surface abundances.
Cataclysmic variables, or CVs, are stably mass-transferring systems with white dwarf (WD) accretors and MS donors. Typically, the progenitor binaries of CVs are post-common-envelope systems (Paczynski et al., 1976; Belczynski et al., 2005; Toonen & Nelemans, 2013; Camacho et al., 2014; Ablimit et al., 2016; Zorotovic & Schreiber, 2022) which have been tightened by magnetic braking and gravitational waves until the onset of Roche lobe overflow (Knigge, 2011).
@@
    the typical accretor mass in CVs is $\simeq~{}0.8M_{\odot}$, greater than the typical observed mass of carbon–oxygen white dwarfs (CO WDs) $\simeq~{}0.6M_{\odot}$, and
@@
    only a small handful of helium(-core) white dwarf (hereafter He WD) accretors have ever been discovered in accreting systems, even though they should be frequent, observable outcomes of conventional binary stellar evolution (Zorotovic et al., 2011; Zorotovic & Schreiber, 2020; Pala et al., 2022).
@@
Figure 1: A summary diagram of the behavior of the He WD–MS merger remnants investigated in this work. On the RGB, merger remnants have cold and slightly more compact cores, which inflates their radii and changes the radiative structures of their cores to which asteroseismology is sensitive (Section 3). Merger remnants which are sufficiently massive enough to reach the tRGB ignite helium in abnormally energetic helium flashes. For massive enough progenitor He WDs, these events can dredge up significant amounts of helium into merger remnants’ envelopes (Section 4), changing their surface abundances, and greatly modifying both their structure and evolution. In our models, such core dredge-up events occur when $M_{\mathrm{WD}}\geq 0.27M_{\odot}$, although the exact threshold depends on other factors such as the cooling age of the He WD (see Appendix A). Less massive remnants become horizontal branch stars (Section 4.5), hot subdwarfs, or He WDs (not explored in this work).
In the past decade, it was realized that the WD mass problem can naturally be solved by an extra, accretor mass-dependent angular momentum loss mechanism (Schreiber et al., 2015). This so-called consequential angular momentum loss would preferentially destabilize the mass transfer of CV-like systems with lower-mass white dwarfs, causing them to quickly merge before they can be observed (Belloni et al., 2018). Physically, this mechanism may be a frictional effect associated with nova events (Shen, 2015; Shen & Quataert, 2022), which are expected to be much longer-duration for lower-mass white dwarfs (Shara et al., 1993) and may cause mergers to occur on the timescale of hours (Shen, 2015), although the subsequent analysis is agnostic to the details.
@@ 2 Stellar models
Figure 2: A summary of the evolution of one possible progenitor system of a He WD–MS merger remnant (left), juxtaposed against our procedure for constructing the stellar model of a merger remnant in MESA (right).
We do this in the following stages (summarized in Figure 2):
@@
    An isolated star with $M=1.2M_{\odot}$ is evolved through the MS and part of the RGB until it attains the desired initial core mass of the remnant, $M_{\mathrm{WD}}$. This initial mass is chosen to be close to typical RG masses from Kepler (e.g., Yu et al., 2018). Because single RGs with $M\lesssim 2M_{\odot}$ all obey the same core mass–luminosity relation, the subsequent analysis is insensitive to this choice.
@@
    To stellar-engineer the merger remnant, we start with a scaffold RG model with core mass $M_{\mathrm{WD}}$. To build the scaffold model, we start with the original RG model from above, and remove envelope mass from the scaffold model via relaxation with $|\dot{M}|=10^{-3}M_{\odot}\,\mathrm{yr}^{-1}$ with nuclear burning disabled, until the RG attains the desired final envelope mass $M_{\mathrm{MS}}$ (physically equal to the mass of the MS star participating in the merger, assuming no mass lost).
    Next, we replace shells with mass coordinates $m\leq M_{\mathrm{WD}}$ with the He WD model, excluding the He WD’s atmosphere (shells where $X\geq 10^{-4}$). Because the modified RG model’s core is now more compact, we recalculate the radial coordinate grid to be consistent with $\mathrm{d}m=4\pi r^{2}\rho\,\mathrm{d}r$. This produces a merger remnant with total mass $M_{\mathrm{tot}}=M_{\mathrm{MS}}+M_{\mathrm{WD}}$.
    For most merger remnant models in this work, we somewhat arbitrarily fix $M_{\mathrm{MS}}$ such that $M=0.8M_{\odot}$, to resemble fairly typical low-mass RGs which may result from CV mergers.
@@
    To ensure numerical convergence, the resulting remnant is evolved for $2\,\mathrm{kyr}$ without burning and gold tolerances disabled. Nuclear reactions and gold tolerances are re-enabled at $2\,\mathrm{kyr}$ and $4\,\mathrm{kyr}$, respectively. Although not initially so, the merger remnant model quickly reaches hydrostatic equilibrium. Evolution through the helium flash involves disabling gold tolerances again, and the models are terminated after helium-burning when the central helium fraction drops to $Y_{c}\leq 10^{-3}$.
@@ 3 Red giant branch
Figure 3: The trajectories on the Hertzsprung–Russell diagram of merger remnant models with varying He WD mass $M_{\mathrm{WD}}$ and total mass $M_{\mathrm{tot}}=M_{\mathrm{MS}}+M_{\mathrm{WD}}=0.80M_{\odot}$. The red dashed line indicates a non-merged model with $M=0.80M_{\odot}$. The top-left and bottom-left panels zoom into the tRGB and CHeB stages, respectively. The red-outlined star symbols on the top-left panel indicate the location of the helium flash. On the bottom-left panel, only evolution $\gtrsim 5\,\mathrm{Myr}$ following the helium flash is shown, and points are distributed $10\,\mathrm{Myr}$ apart. We have excised times near the helium flash, when the envelope sometimes becomes hydrodynamical.
@@ 3.2 Asteroseismic signatures on the red giant branch
defined here specifically for the dipole ($\ell=1$) modes. This quantity probes the radiative core of the star. The integral in Equation 4 is taken over $\mathcal{R}$, the g-mode cavity of the RG, i.e., the region where $2\pi\nu_{\mathrm{max}}<N$, where $N$ is the Brunt–Väisälä frequency. At present, $\Delta\Pi$ has been measured for a few thousand RGs (e.g., Vrard et al., 2016), observationally constraining the structures of their radiative cores.
@@
However, in a short time $<\tau_{\mathrm{th,core}}$, electronic thermal conduction (approximately obeying the nonlinear law in Equation 18) sets up a temperature gradient $\nabla$. Curiously, we find that this temperature gradient happens to mostly cancel out the factor $\sqrt{T_{\mathrm{shell}}/T}$ (see Equation 5) to cause the $N$ profiles of the single and merger remnant models to be similar (top center panels of Figure 4). Equality between the Brunt–Väisälä frequency profiles occurs long before the interior of the core has thermalized to the shell temperature. Therefore, in the relatively short time that it takes heat conduction to set up a temperature gradient throughout the core, the lower temperature of the merger remnant core no longer works to increase $\Delta\Pi$. The reason for this “coincidental” cancellation remains mysterious, but may be related to some properties of long-lived pre-thermalized solutions of nonlinear heat diffusion. These solutions (called “intermediate asymptotics”; Barenblatt, 1996) exhibit self-similar behavior in many nonlinear heat diffusion problems (often in the context of diffusion in a porous medium), and have been studied extensively in applied mathematics (by, e.g., Witelski & Bernoff, 1998; Galaktionov et al., 2004; Hayek, 2014).
Figure 4: The Brunt–Väisälä frequency $N$ (top) and temperature $T$ (bottom) shown for the $M_{\mathrm{MS}}=0.55M_{\odot}$, $M_{\mathrm{WD}}=0.25M_{\odot}$ merger remnant (solid gray-blue) and non-merged model of $M=0.80M_{\odot}$ (red dashed), at various stages of evolution. The two models are compared at equal core masses $M_{c}$. The shaded region denotes helium deposited onto the helium core of the merger remnant model by hydrogen burning during its evolution.
@@
The g-mode period spacing $\Delta\Pi$ is only observable for remnants sufficiently low on the RGB (i.e., below the RGB bump, which is the “hook” feature in Figure 5). For merger remnants in this regime, departure from the degenerate sequence is most prominent during the initial contraction phase after the formation of the remnant, although RGs which have passed this phase still lie above the degenerate sequence for an additional $\sim 20\,\mathrm{Myr}$. Hence, this particular asteroseismic diagnosis may only be possible for sufficiently recent mergers involving sufficiently low-mass He WDs.
Figure 5: An asteroseismic spacing diagram ($\Delta\Pi$ versus $\Delta\nu$) showing merger remnant evolutionary sequences and a non-merged model (red dashed) all with $M=0.80M_{\odot}$. The top-right panel shows the entire parameter space, and the left and bottom panels zoom in on CHeB stars and the RGB, respectively. Roughly speaking, both RGB and CHeB stars evolve from right to left on this diagram (i.e., towards decreasing $\Delta\nu$, or increasing $R$). The black points are observations taken from the catalog of Vrard et al. (2016) with stars flagged for possible aliases removed. In the top-left panel, colored triangles denote the last $5\,\mathrm{Myr}$ of the CHeB phase. The blue squares denote $8$ stars from the very low-mass sample of Li et al. (2022) which appear in the Vrard catalog.
@@ 3.3 Overbright tip of the RGB
Figure 6: Top: The maximum helium-burning luminosities $L_{\mathrm{He,flash}}$ attained during the helium flash in our merger remnant models with total mass $M_{\mathrm{tot}}=M_{\mathrm{MS}}+M_{\mathrm{WD}}=0.80M_{\odot}$. Bottom: For the merger remnant models, we show core masses $M_{\mathrm{c@flash}}$ at the helium flash (purple circles), mass coordinates $M_{\mathrm{flash}}$ where the helium flash begins (magenta stars), and initial (turquoise triangles) and final (green squares) core masses $M_{\mathrm{CHeB,ci}}$ and $M_{\mathrm{CHeB,cf}}$ during the CHeB phase. In both panels, dashed lines denote values for the $M=0.80M_{\odot}$ non-merged model.
@@ 4 Core helium-burning phase
In our models, the maximum helium-burning luminosity attained by a merger remnant is roughly a linear function of $M_{\mathrm{WD}}$, with the most extreme model attaining a factor of $\approx\!5$ higher helium-burning luminosity than attained by a non-merged star (see the top panel of Figure 6). The energy production rate can be a whopping $\approx\!10^{10}L_{\odot}$ at the peak of the helium flash. Since mergers involving more massive WDs than those on our model grid are possible, even larger helium-burning luminosities may occur in nature. The most vigorous helium flash in our merger models also burns a larger fraction of the core into carbon, up to $\simeq 7\%$ by mass. We present the post-helium flash composition profiles of several of our models in Appendix D.
@@ 4.3 Asteroseismic mixed-mode coupling
Figure 11: Selected surface abundances $\mathrm{A}(\mathrm{X})$ (defined in Equation 11) during the CHeB phase. Red points denote the $M=0.80M_{\odot}$ non-merged model, and starred points denote post-dredge-up merger remnants. For ${}^{7}\mathrm{Li}$, filled symbols denote lithium enrichment due to the helium flash only (described in Section 4.4), whereas hollow symbols denote lithium enrichment due to both the helium flash and merger itself (which may be inaccurately modelled). Note that the vast majority of CHeB stars have $A(\mathrm{Li})\lesssim 1.5$ (Kumar et al., 2020), and the solar abundance is $A(\mathrm{Li})\approx 1.05$ (Asplund et al., 2009).
@@ 4.4 Abundance anomalies
We find that post-dredge-up remnants exhibit significant 12C surface enrichment, up to $\simeq 1\,\mathrm{dex}$ relative to hydrogen. This enhancement arises from dredge-up of core material which has been partially fused during the helium flash into carbon. Post-dredge-up remnants thus possess increased values of $\mathrm{C}/\mathrm{Fe}$ and decreased values of ${}^{13}\mathrm{C}/{}^{12}\mathrm{C}$. Additionally, our post-dredge-up remnant models also possess significant surface enhancements (up to $\sim 1.5\,\mathrm{dex}$ relative to hydrogen) in ${}^{18}\mathrm{O}$ and ${}^{22}\mathrm{Ne}$, which are created by successive $\alpha$ captures of ${}^{14}\mathrm{N}$ (e.g., Clayton, 2003) during the helium flash. As the surface abundance of ${}^{16}\mathrm{O}$ in all of our merger remnant models is almost identical to that of our non-merged model, the surface abundance ratio ${}^{16}\mathrm{O}/^{18}\mathrm{O}$ is also decreased by up to $\sim 1.5\,\mathrm{dex}$ in post-dredge-up remnants. While not included in our reaction network, it is also probable that a significant amount of ${}^{26}\mathrm{Mg}$ is formed through $\alpha$ capture of ${}^{22}\mathrm{Ne}$ in the abnormally hot helium flash (as in Shen et al., 2023). The surface abundance of ${}^{26}\mathrm{Mg}$ is therefore also likely to be enhanced.
The lithium abundance is also important, because it can be created by the burning of 3He or destroyed by burning at temperatures comparable to those required for hydrogen burning. Lithium-rich giants make up about $3\%$ of all CHeB stars (Kumar et al., 2020), and have previously been suggested to have formed via binary interactions or mergers (e.g., Zhang & Jeffery, 2013; Casey et al., 2019). Our merger remnant models with high $M_{\mathrm{WD}}$ become highly lithium-rich very soon after merger, owing to a brief dredge-up event which occurs when hydrogen burning is first turned on (similar to the Cameron–Fowler mechanism; Cameron & Fowler, 1971). While such a merger-era lithium enhancement seems plausible (see, e.g., observations of lithium enrichment in luminous red novae; Kamiński et al., 2023), our treatment of the merger process itself is highly artificial, and the evolution of the remnant is unreliable for post-merger ages younger than a thermal time, $\tau_{\mathrm{th,env}}$. Modification to the surface abundances of other species during this (possibly artificial) dredge-up event are minor, since they primordially occur in much higher abundance than lithium.